OpenStudy (anonymous):

How would I convert Θ=2 into degrees, minutes, and seconds?

OpenStudy (mertsj):

The conversion factor is 180 degrees = pi radians

OpenStudy (anonymous):

\[2/\pi= \pi/180 \] Is that how it should look? Thank you, by the way.

OpenStudy (mertsj):

\[2 radians \times \frac{180 ^{o}}{\pi radians}\]

OpenStudy (anonymous):

Okay, so I have \[360/\pi\], can I multiply again by \[\pi/2\] to get rid of the pi symbol?

OpenStudy (anonymous):

Nevermind! Sorry, I just worked that out and it's completely wrong.

OpenStudy (mertsj):

Just divide by 3.14159